Assume that we want to construct a confidence interval. Do one of the​ following, as​ appropriate: (a) find the critical value t Subscript alpha divided by 2​, ​(b) find the critical value z Subscript alpha divided by 2​, or​ (c) state that neither the normal distribution nor the t distribution applies. Here are summary statistics for randomly selected weights of newborn​ girls: nequals161​, x overbarequals29.9 ​hg, sequals6.5 hg. The confidence level is 95​%.
Answer #1

solution

n = 161

Degrees of freedom = df = n - 1 = 161- 1 = 160

At 95% confidence level the t is ,

= 1 - 95% = 1 - 0.95 = 0.05

  / 2= 0.05 / 2 = 0.025

t /2,df = t0.025,160 = 1.975 ( using student t table)

in two decimal answer=1.97